# Intelligent Tutors for Adult Learners: An Analysis of Needs and Challenges

### v2

@@ 3.5. Hinting Strategy
Each tutor has a dedicated hint box at the bottom of the problem interface, which provides users with the ability to request help. Multiple layers of hints were implemented for each problem type, allowing users to access up to three tiers of assistance (as seen in Figure 1, section D). In the initial hint, the specific step under consideration is highlighted and the user is prompted with a message encouraging them to tackle the particular step. At the second hint level, an explanatory message is delivered in the hint box, offering guidance on how to approach solving the step. In the final hint tier, the solution to the step is revealed, effectively providing the learner with a worked example.
Providing the solution in the final hint tier serves several purposes. Research shows that worked examples are very effective for learning, especially for beginners, as they reduce mental effort and help students learn problem-solving skills (Renkl, 2002). Showing the correct solution ensures that learners see the right approach, even if they cannot figure it out on their own. This helps reinforce key ideas and allows learners to compare their own attempts with the correct method, improving understanding and memory (VanLehn, 1996). Allowing learners to enter the revealed solution also keeps them actively involved, rather than just passively looking at the answer. Every time a hint is used, it is logged in the Apprentice Tutors database for further analysis of how users interact with hints.
3.6. Adaptive Problem Selection
@@
The Apprentice Tutors platform provides a personalized tutoring experience through the adaptive problem selection option (as seen in Figure 1, Section A). While each Apprentice tutor has this adaptive selection tab, this is optional, and learners may opt to click on a problem type directly. The adaptive problem selection option personalizes the student’s learning experience by providing them with problems that exercise skills they have not yet mastered. Actions such as entering a tutor step or requesting a hint are logged within the tutoring database. Apprentice Tutors uses these logs to track the progression of students’ skills. Each of these logged steps is referred to as a tutor transaction. Each tutor transaction references skills within the tutor’s problem type, represented by knowledge components (also referred to as KCs). The hypothesized KCs are defined as acquired units of cognitive function that can be inferred from the performance of a set of related tasks (Koedinger et al., 2012). Bayesian Knowledge Tracing (BKT) (Corbett and Anderson, 1994) is used to assess the student’s mastery of each skill. BKT is an approach that models a student’s learning using a hidden Markov model. It is commonly used in intelligent tutoring systems (Yudelson et al., 2013). The idea behind BKT is that a student’s competency is not fixed; as a student interacts with a problem, their competency improves. BKT assumes that initially the student may not know a concept but can achieve mastery of the concept over time with practice. Based on BKT’s estimates, adaptive problem selection randomly selects problems for students to practice that exercise at least one unmastered skill.
3.7. Performance Visualization Dashboards
@@
Visualizations of student trajectories are increasingly used to comprehend student activity and their navigation through course content (Goulden et al., 2019; Lundqvist and Warburton, 2019). To facilitate the visualization of skill progression, I developed a dashboard for students to view their mastery on each skill. This dashboard displays the mastery level of each KC, broken down by the specific problems in which these KCs are encountered (as seen in Figure 1, part E). In addition, Guo et al. (Guo et al., 2024) developed visual analytics dashboard that shows detailed provenance data across multiple coordinated views (Guo et al., 2024).
Visualization dashboards are used for several key purposes. For students, these dashboards provide immediate feedback on their progress, helping them identify areas of strength and weakness, which can promote self-regulated learning (Kanfer et al., 2022). For teachers, they offer information on the performance of individual and class-level learners, allowing targeted interventions and more informed instructional decisions (Vanlehn, 2006).
4. Study 1: Tutor Adoption Across Deployments
To investigate the challenges that prevent adult learners from adopting intelligent tutors, I deployed the Apprentice Tutors as an optional supplement in several sections of a College Algebra. A substantial body of research has consistently demonstrated improvements in user competency through practice with intelligent tutoring systems (Pane, 1967; Koedinger et al., 1997). However, this research does not assess usage and adoption rates, particularly in contexts where learners can choose whether to use these systems and there are no extrinsic motivators (e.g., payment or grades). In addition, current work mainly focuses on the effects of the deployment of intelligent tutors in K–12 educational settings (Koedinger et al., 1997; Heffernan and Heffernan, 2014; Aleven et al., 2006). To extend these findings, my tutor deployment studies focus on adult learners, for whom participation is not compulsory. By examining patterns of tutor adoption and usage over time, along with the specific needs of adult learners, valuable insights can be provided to educational technology developers seeking to achieve a sustained impact.
@@ 4.1. Methods
I made Apprentice Tutors for four college algebra topics available at several state technical college classes spanning from September 2021 to June 2022. The institution, which focuses on technical 2-year degree programs, primarily serves non-traditional adult learners returning to school to acquire new skills in support of their career goals.
To address limitations in existing research, this study focuses on adult learners as a key user population for intelligent tutoring systems. In this context, “adult learners” are defined students over the age of 18, with an emphasis on non-traditional learners that have already spent time in the workforce. This group often includes older individuals who have already spent time in the workforce and are returning to education to acquire new skills or advance their careers. These learners often balance their studies with personal and professional responsibilities, making their needs and motivations distinct from those of younger, traditional college-aged students. Adult learners present unique challenges and opportunities due to their distinct motivations, cognitive developmental differences, and the voluntary nature of their involvement with educational technology.
This study examines how sociotechnical factors influence tutor adoption and usage in self-directed learning contexts, focusing specifically on adult learners. By analyzing their lived experiences and unique needs, we aim to improve the design of intelligent tutors to better support adult learner’s educational goals and skill acquisition. The ultimate goal is to enhance learning outcomes for adults, enabling them to succeed in both personal and professional development. Lifelong learning and skill development are increasingly critical in today’s rapidly evolving workforce, where continuous education is essential for maintaining employability and competitiveness (Bell and Blanchflower, 2017).
During the deployment, students accessed the tutors through the Blackboard Learn system. The use of the tutors was entirely voluntary, and no course credit or payment was offered for their use. Before and during the deployment, we engaged teachers through email communications and presentations at their monthly department meetings, providing information about the tutors, their features, and potential benefits for students. Before starting this study, its protocol was reviewed and approved by my institutional review board. Upon first accessing a tutor, students were asked to consent to participate in the study. Those who did not consent could still use the tutor, but their data was neither collected nor analyzed. For consenting students, data was collected on usage, including access times, tutor interactions, and student-course correlations.
@@ 4.2.1. Usage and Adoption
During the deployment period, 3,510 students were granted access to the Apprentice Tutors. Figure 2 illustrates the usage and retention funnel throughout the first four academic quarters of deployment. Adoption was defined as the percentage of students who clicked on the tutors at least once, resulting in an overall adoption rate of approximately 14.8%. In particular, as the number of classes that use the system increased from 4 to 58, the rate of initial clicks decreased from 62. 84% to 13. 78%. A term-by-term breakdown of tutor deployment and adoption detlays is stated in Table 1.
To evaluate tutor adoption, we assessed the percentage of students who accessed the tutors and solved more than one problem. This metric decreased slightly over time, from 17% in the first term to 14% in the fourth term. Engagement was highest during the second term, when 22% of students who initially accessed the tutors proceeded to solve more than one problem.
@@ 4.3. Discussion
5. Study 2: Qualitative Investigation of Challenges
The deployment of Apprentice Tutors showed a positive trend in user retention, where nearly half of the users who solved at least one problem continued to solve at least five or more unique problems. Despite this, the overall adoption rate of these tutoring systems remained at 14.8%. My analyses suggest an inverse relationship between the deployment scale and adoption rates. A plausible reason for the lower adoption of tutors over time may have been the decrease in direct teacher support as the tutor deployment scaled. This support included an email nudge approach to remind teachers to further remind their students that tutors were available as a resource to use.
@@ 5.1.1. Recruitment
To investigate my research questions, we conducted five semi-structured focus group sessions with both students (who used the tutors) and teachers (who agreed to tutors being deployed in their classrooms). We recruited participants through snowball sampling, through direct e-email, and newsletter postings. Throughout the five sessions, we collected data from fifty-three participants, which consisted of forty-two students and eleven teachers. Participation in these focus groups was optional and a 15 dollar gift card was available to participants after the completion of the session. For the purposes of these focus groups, age and demographics were not used in the filtering or selection of the participants. All participants were located in the United States and all focus groups were conducted virtually. All virtual focus groups were about 60 minutes.
@@ 5.1.3. Data Collection
The prompts presented were open-ended and semi-flexible, where participants were given the opportunity to reflect more broadly on their experiences through other academic experiences.
Intelligent tutor key themes Participant Quotes from Focus Groups
@@ 5.2. Results
The focus group data was analyzed to identify key adult learners’ needs that AI-powered tutoring systems could address. A bottom-up analytical approach was implemented within Miro, where raw data was mapped onto color-coded digital post-it notes. Through iterative analysis, several emergent themes were identified, which are summarized in Table 2. This table presents the main theme categories, specific themes, and exemplar user needs within each theme.
@@
We received recurring feedback on tutor requirements through focus groups with teachers and students. The data suggested that adoption might improve if the tutors were: (a) easy to understand, (b) easy to navigate, and (c) user-friendly. One teacher described this perspective as follows:
“Now the concept is great. What I was thinking when I first heard of this is, you know I don’t delve deeply into simplifying radicals because I don’t have the time. If I had somewhere where they could go practice on simplifying radicals, that would be great. But if it was one that was easy to understand and easy to get through and user friendly, Amen.”
@@
  • AI Support Feelings: Concerns were raised that tutors could reduce the total amount of “human” or “human-like” interaction (“It takes out the human factor… the AI isn’t gonna understand that there might be some barriers…”), but users also valued “superhuman” capabilities beyond what a single instructor can provide (“If they can get more practice on that topic… maybe they can get more practice on that topic if they’re struggling…”).
@@ 5.3. Discussion
The focus group sessions and subsequent data analysis revealed a complex interplay of factors that can hinder the wider adoption and sustained use of tutoring systems among adult learners. Although the results section highlighted several positive aspects and demonstrated the potential of these systems, a number of additional insights surfaced that point to challenges in real-world implementation.
@@
It was also observed that, beyond simple measures of tutor effectiveness, other sociotechnical factors strongly influenced adoption. These factors include ease of use, technological infrastructure, scheduling constraints, and alignment with curriculum objectives. Adult learners, in particular, can face additional barriers such as limited time, competing responsibilities, and varying levels of digital fluency. These elements underscore the importance of holistic approaches to designing and deploying AI tutors that address not only instructional goals but also practical considerations of user contexts.
@@ 6. Limitation and Future Works
In addition, future work should expand tutor coverage to encompass a broader range of course material, and model tracing will be integrated alongside knowledge tracing and learning analytics visualizations. These enhancements will provide teachers (and students) with more robust tools to track progress and better understand the problem-solving steps. Future research will involve redeploying these improved tutors and examining their impact over time through longitudinal studies across multiple deployments. Such investigations will facilitate a deeper understanding of how to develop scalable pedagogical technologies that effectively support adult learning.
@@ 7. Conclusion
Intelligent tutoring systems have demonstrated improvements in users’ learning outcomes, offering an effective and more cost-efficient alternative to human tutoring (VanLEHN, 2011). There is an unexplored potential in understanding how intelligent tutors can impact adult learners and how this population might adopt educational technologies. Further research is needed to investigate these opportunities and address this underexplored area. The Apprentice Tutors were deployed as a case study to explore the specific needs of adult learners who engage with intelligent tutoring systems. The preliminary research yielded quantitative tutor usage and user retention after several problems were completed. Despite these positive results, a gradual decline in the adoption of these tutors was observed over time. This observation led to a subsequent study that aimed to elicit teacher and student experiences with and perceptions of intelligent tutors in adult learning contexts. In this second phase, focus groups were conducted to collect qualitative insights from both students and teachers who interacted with the tutors. Analysis of this data through thematic analysis revealed several themes that informed a strategic roadmap for enhancing the Apprentice Tutors platform. Furthermore, these insights were synthesized into general recommendations for developers and instructional designers, aimed at improving the engagement and adoption rates of intelligent tutoring systems more broadly. The goal of this research is to contribute to the widespread implementation of intelligent tutors, ensuring that these advanced educational tools are accessible and equitable to adult learners seeking opportunities for educational growth.

### v3

@@ 3.1.4. Hinting Strategy
Each tutor has a dedicated hint box at the bottom of the problem interface, which provides users with the ability to request help. Multiple layers of hints were implemented for each problem type, allowing users to access up to three tiers of assistance (as seen in Figure 1, section D). In the initial hint, the specific step under consideration is highlighted and the user is prompted with a message encouraging them to tackle the particular step. At the second hint level, an explanatory message is delivered in the hint box, offering guidance on how to approach solving the step. In the final hint tier, the solution to the step is revealed, effectively providing the learner with a worked example. This hinting strategy is inspired by Anderson et al. (1995).
3.1.5. Adaptive Problem Selection
@@
The Apprentice Tutors platform provides a personalized tutoring experience through the adaptive problem selection option (as seen in Figure 1, Section A). While each Apprentice tutor has this adaptive selection tab, this is optional, and learners may opt to click on a problem type directly. The adaptive problem selection option personalizes the student’s learning experience by providing them with problems that exercise skills they have not yet mastered. Actions such as entering a tutor step or requesting a hint are logged within the tutoring database. Bayesian Knowledge Tracing (BKT) (Corbett and Anderson, 1994) is used to assess the student’s mastery of each skill. BKT is an approach that models the student’s learning using a hidden Markov model. It is commonly used in intelligent tutoring systems (Yudelson et al., 2013).
3.1.6. Performance Visualization Dashboards
@@
Visualizations of student trajectories are increasingly used to comprehend student activity and their navigation through course content (Goulden et al., 2019; Lundqvist and Warburton, 2019). To facilitate the visualization of skill progression, we developed a dashboard for students to view their mastery on each skill. This dashboard displays the mastery level of each KC, broken down by the specific problems in which these KCs are encountered (as seen in Figure 1, part E). In addition, Guo et al. (Guo et al., 2024) developed visual analytics dashboard for instructors that shows detailed provenance data across multiple coordinated views (Guo et al., 2024).
@@ 3.2. Methods
We developed tutors for four college algebra topics available at several state technical college classes spanning from September 2021 to June 2022. The institution, which focuses on technical 2-year degree programs, primarily serves non-traditional adult learners returning to school to acquire new skills in support of their career goals. Adult learners present unique challenges and opportunities due to their distinct motivations, cognitive developmental differences, and the voluntary nature of their involvement with educational technology. This study examines how sociotechnical factors influence tutor adoption and usage in self-directed learning contexts, focusing specifically on adult learners. By analyzing their lived experiences and unique needs, we aim to improve the design of intelligent tutors to better support adult learner’s educational goals and skill acquisition. The ultimate goal is to enhance learning outcomes for adults, enabling them to succeed in both personal and professional development.
During the deployment, students accessed the tutors through the Blackboard Learn system. The use of the tutors was entirely voluntary, and no course credit or payment was offered for their use. Before and during the deployment, we engaged teachers through email communications and presentations at their monthly department meetings, providing information about the tutors, their features, and potential benefits for students. Before starting this study, its protocol was reviewed and approved by our institutional review board. Upon first accessing a tutor, students were asked to consent to participate in the study. Those who did not consent could not use the tutors. For consenting students, data was collected on usage, including access times, tutor interactions, and student-course correlations.
@@ 3.3.1. Usage and Adoption
During the deployment period, 3,510 students were granted access to the Apprentice Tutors. Figure 2 illustrates the usage and retention funnel throughout the first four academic quarters of deployment. Adoption was defined as the percentage of students who clicked on the tutors at least once, resulting in an overall adoption rate of approximately 14.8%. In particular, as the number of classes that use the system increased from 4 to 58, the rate of initial clicks decreased from 62.84% to 13.78%. A term-by-term breakdown of tutor deployment and adoption metrics are stated in Table 1.
To evaluate tutor adoption, we assessed the percentage of students who accessed the tutors and solved more than one problem. This metric decreased slightly over time, from 17% in the first term to 14% in the fourth term. The participation was highest during the second term, when 22% of the students who initially accessed the tutors proceeded to solve more than one problem.
@@ 3.4. Discussion
4. Study 2: Investigation of Challenges
The deployment of Apprentice Tutors showed a positive trend in user retention, where nearly half of the users who solved at least one problem continued to solve at least five or more unique problems. Despite this, the overall adoption rate of these tutoring systems remained at 14.8%. Our analyses suggest an inverse relationship between the deployment scale and adoption rates. A plausible reason for the lower adoption of tutors over time may have been the decrease in direct teacher support as the tutor deployment scaled. This support included an email nudge approach to remind teachers to further remind their students that tutors were available as a resource to use.
@@ 4.1.1. Recruitment
To investigate our research questions, we conducted five semi-structured focus group sessions with both students (who used the tutors) and teachers (who agreed to tutors being deployed in their classrooms). We recruited participants through snowball sampling, through direct e-email, and newsletter postings. Throughout the five sessions, we collected data from fifty-three participants, which consisted of forty-two students and eleven teachers. Participation in these focus groups was optional and a 15 dollar gift card was available to participants after the completion of the session. For the purposes of these focus groups, age and demographics were not used in the filtering or selection of the participants. All participants were located in the United States and all focus groups were conducted virtually. All virtual focus groups were about 60 minutes.
@@ 4.1.3. Data Collection
The prompts presented were open-ended and semi-flexible, where participants were given the opportunity to reflect more broadly on their experiences through other academic contexts.
Key Intelligent Tutor Themes Participant Quotes from Focus Groups
@@ 4.2. Results
We identified key key adult learners’ needs that AI-powered tutoring systems could address through analyzing the focus group data. To do so, we conducted a bottom-up analytical approach within Miro, where raw data was mapped onto color-coded digital post-it notes. Through iterative analysis, we identified several emergent themes, which are summarized in Table 2. This table presents the main theme categories, specific themes, and exemplar user needs within each theme.
@@
We received recurring feedback on tutor requirements through focus groups with teachers and students. The data suggested that adoption might improve if the tutors were: (a) easy to understand, (b) easy to navigate, and (c) user-friendly. One teacher described this perspective as follows: “Now the concept is great. What I was thinking when I first heard of this is, you know I don’t delve deeply into simplifying radicals because I don’t have the time. If I had somewhere where they could go practice on simplifying radicals, that would be great. But if it was one that was easy to understand and easy to get through and user friendly, Amen.”
@@
  • AI Support Perceptions: Concerns were raised that tutors could reduce the total amount of “human” or “human-like” interaction (“It takes out the human factor… the AI isn’t gonna understand that there might be some barriers…”), but users also valued “superhuman” capabilities beyond what a single instructor can provide (“If they can get more practice on that topic… maybe they can get more practice on that topic if they’re struggling…”).
@@ 4.3. Discussion
The focus group sessions and subsequent data analysis revealed a complex interplay of factors that can hinder the wider adoption of tutoring systems among adult learners. Although the results section highlighted several positive aspects and demonstrated the potential of these systems, a number of additional insights surfaced that point to challenges in real-world implementation.
@@
We found that sociotechnical factors played an important role in the adoption of tutors beyond just effectiveness. These factors include ease of use, technological infrastructure, scheduling constraints, and alignment with curriculum objectives. Adult learners, in particular, can face additional barriers such as limited time, competing responsibilities, and varying levels of digital fluency. These elements underscore the importance of holistic approaches to designing and deploying AI tutors that address not only instructional goals but also practical considerations of user contexts.
@@ 5. Limitation and Future Works
In addition, future work should expand tutor coverage to encompass a broader range of course material, and model tracing will be integrated alongside knowledge tracing and learning analytics visualizations. These enhancements will provide users with more robust tools to track progress and better understand the problem-solving steps. Future research will involve redeploying these improved tutors and examining their impact over time through longitudinal studies across multiple deployments. Such investigations will facilitate a deeper understanding of how to develop scalable pedagogical technologies that effectively support adult learning.
@@ 6. Conclusion
There is an unexplored potential in understanding how intelligent tutors can impact adult learners and whether population would adopt educational technologies. The Apprentice Tutors were developed and deployed as a case study to explore the specific needs of adult learners who engage with intelligent tutoring systems. Our preliminary work yielded quantitative tutor usage and user retention after several problems were completed. Despite these positive results, a gradual decline in the adoption of these tutors was observed over time. This observation led to a subsequent study that aimed to elicit teacher and student experiences with and perceptions of intelligent tutors in adult learning contexts. In this second study, we conducted focus groups to collect qualitative insights from both students and teachers who interacted with the tutors. Analysis of this data through thematic analysis revealed several themes that informed a strategic roadmap for enhancing the Apprentice Tutors platform. Furthermore, these insights were synthesized into general recommendations for developers and instructional designers, aimed at improving the engagement and adoption rates of intelligent tutoring systems more broadly. The goal of this research is to contribute to the widespread implementation of intelligent tutors, ensuring that these advanced educational tools are accessible and equitable to adult learners seeking opportunities for educational growth.